OpenStudy (anonymous):

The function H(t) = −16t^2 + 60t + 95 shows the height H(t), in feet, of a projectile after t seconds. A second object moves in the air along a path represented by g(t) = 20 + 38.7t, where g(t) is the height, in feet, of the object from the ground at time t seconds.

OpenStudy (anonymous):

Part A: Create a table using integers 1 through 4 for the 2 functions. Between what 2 seconds is the solution to H(t) = g(t) located? How do you know? (6 points) Part B: Explain what the solution from Part A means in the context of the problem. (4 points) @e.mccormick @welshfella @ali2x2

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@FEARLESS_JOCEY @ganeshie8

OpenStudy (anonymous):

@nincompoop @love10129151 @SolomonZelman @mathstudent55

OpenStudy (anonymous):

please?

OpenStudy (e.mccormick):

Well, first, did you chart the points they are taling about?

OpenStudy (anonymous):

no? @e.mccormick

OpenStudy (anonymous):

I am totally confuzzled

OpenStudy (anonymous):

I graphed it, but idk where to go from there

OpenStudy (e.mccormick):

Well, you need to chart out the points. it will be something like: \(\begin{matrix} t= & 1 & 2 & 3 & 4\\ H(t) & & & & \\ g(t) & & & & \end{matrix} \)

OpenStudy (anonymous):

okayyy

OpenStudy (anonymous):

dinner soon- can we speed this up?

OpenStudy (e.mccormick):

So lets look at 1. You put it in each equation. \(\begin{array}{|c|c|c|c|}\hline t= & 1 & 2 & 3 & 4\\ \hline H(t) & 139 & & & \\ \hline g(t) & 58.7 & & & \hline \\ \hline \end{array}\) And you just fill it in.

OpenStudy (e.mccormick):

All it is saying is that at t=1, H(t) = 139 and g(t) = 58.7. You just plug the number in to each equation to get these answers.

OpenStudy (anonymous):

h(t), t=2, h(t)=151

OpenStudy (e.mccormick):

Yep.

OpenStudy (anonymous):

g(t), t=2, g(t)=97.4

OpenStudy (anonymous):

i just read all this and im sorry but i have no clue

OpenStudy (anonymous):

g(t)=136.1, when t=3

OpenStudy (anonymous):

h(t), when t=3, h(t)=131

OpenStudy (anonymous):

good so far?

OpenStudy (anonymous):

h(t)=79, when t=4

OpenStudy (e.mccormick):

\(\begin{array}{|r|c|c|c|}\hline t= & 1 & 2 & 3 & 4\\ \hline H(t)= & 139 & 151 & 131 & \\ \hline g(t) =& 58.7 & 97.4 & 136.1 & \hline \\ \hline \end{array}\) Yah, fine.

OpenStudy (e.mccormick):

\(\begin{array}{|r|c|c|c|}\hline t= & 1 & 2 & 3 & 4\\ \hline H(t)= & 139 & 151 & 131 & 79 \\ \hline g(t) =& 58.7 & 97.4 & 136.1 & 174.8 \\ \hline \end{array}\)

OpenStudy (e.mccormick):

OK... so, there is the chart. So the question about the chart is " Between what 2 seconds is the solution to H(t) = g(t) located? How do you know? "

OpenStudy (e.mccormick):

If you look at the chart and the equations, you know that H(t) is a curve and g(t) is a line that intersect somewhere. Now H(t) obviously goes up from 1 to 2, but some time between 1 and 3 it peaked and is going down. With all that info, can you see where they might cross?

OpenStudy (e.mccormick):

I seem to have lost you, so whenever you get back, look at this: https://www.desmos.com/calculator/dam9lkolsw